3sinx - 4cosx = 0 then what is the value of cotx is ????

Grade:10

2 Answers

Nishant Vora IIT Patna
7 years ago
dear student,

3sinx = 4cosx

hence , cosx/sinx = ¾
so cot x = ¾

Thanks
RAMCHANDRARAO
7 years ago
 3sinx-4cosx=0 that implies 3sinx=4cosx that implies 3=4cotx
therefore cotx=3/4 then tanx=4/3 in this way we can solve this type of problems